OpenStudy (anonymous):

Help please, anyone? Layla has a jar filled with 108 coins. If the jar contains only quarters and nickels and has a value of $15.80, how many nickels are in the jar? 52 53 55 56

OpenStudy (anonymous):

I think you have to do system of equations here. Since you know there are two types of coins and there are a total of 108 coins. Then your first equation would be\[x+y=108\]Next, you know a quarter is 25 cents and a nickel is 5 cents. Total is $15.80.\[(.25)x+(.05)y=15.80\]Then do substitution

OpenStudy (anonymous):

what substitution?

OpenStudy (anonymous):

Get either x or y by itself from the first equation and then plug x or y into the second equation to find the x or y. It depends if you chose x to get by itself or y by itself.

OpenStudy (anonymous):

Agh, I'm still confused, could you please show me the work step by step?

OpenStudy (anonymous):

Ok so lets say I want to get x by itself from the first equation \[x+y=108 => x=108-y\]Then plug x into the second equation\[(.25)(108-y)+(.05)y=15.80\]Then distribute and solve for y \[27-0.25y+0.05y=15.80 => -0.20y=-11.20 => y=56\] since I chose y to be nickels (from the second equation). That means there are 56 nickels.

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

Yep no problem